OpenStudy (babyslapmafro):

Can someone demonstrate or explain how to differentiate the following function by using the product rule? f(x)=(2x^2-3)(x^2+5)

OpenStudy (babyslapmafro):

I differentiated by multiplying first and got 8x^3+14x

OpenStudy (babyslapmafro):

But I am not sure how to use the product rule.

OpenStudy (anonymous):

f(x)=u(x)+v(x) f'(x)=u'(x)v(x)+u(x)v'(x) u(x)=2x^2-3 u'(x)=4x v(x)=x^2+5 v'(x)=2x f'(x)=4x*(x^2+5)+(2x^2-3)*2x f'(x)= ...............